Home‎ > ‎Reports‎ > ‎Crew103‎ > ‎

day10


Crew 103 Log Book for April 18, 2011

Photos of the day





Commander's Check-In Report

Anne-Sophie Kalbfleisch Reporting

Crew Physical Status: everybody is fine

Time Departed/Returned from EVA: 16h-18h

Brief Narrative of Field Mission Results:

- EVA: 3 out of the 4 soil samples have been brought back

- The analysis of our algae's is continuing

- The basic structure of our hot-air balloon is constructed

EVA Data/Interpretations:

/

Engineering/Hab Maintenance:

Nothing to mention

Plans for Tomorrow:

We will continue our hot-air balloon experience, take other measures

of the optical density of our algae's (we have sent an email with some

concerns about our results to the University of Liège which with we

are working), we will try to make an internet relay for our max rover

so that we could go farther with him and we will start the humidity

analysis of our soil samples.

Support Requested:

Nothing requested

Miscellaneous:

/


Commander's Report

Anne-Sophie Kalbfleisch Reporting

Today we spent a lot of time filling our psychological reports, for our Japanese psychologist as well as for our Belgian psychologist, which are following us.

T. Windels and S. Viroux went on EVA because the team of yesterday had forgotten the little plastic bags to take the soil samples. They had to find this places back thanks to the GPS coordinates noticed by the previous team.

During the afternoon we started a new exciting experience. Crew 94, which was here just a year ago, had started an experience with some balloons filled with helium in order to take some aerial photographs of the surroundings. But this experience failed and they never succeed in sending the balloons in the air. Now we plan to do the same with a self-made hot-air balloon. Today we just constructed the basic structure and tomorrow, when the wind will be less strong, we will try to fill it with hot air outside, thanks to our Bunsen burner. More details about it in the science report.

I forgot to mention it in the previous reports : each two days T.

Windels and I are doing one hour of gym because some members of crew 94 (JUMP) told us that they felt kind of lazy after one week doing only experiments and that they felt a lot better the second week after doing some gym. Mens sana in corpore sano :)



Engineering Report

Thibault Windels Reporting

Generator/Electricity:

Honey (Diesel Generator) (Used? Yes/No): yes

Wendy (Gas Generator) (Used? Yes/No): no

Battery Charge Level: 10,3 volts


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0,49

Propane Tank (Full/0.75/0.5/0.25/Empty): 0,38

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 6 galons approximately

Notes/Comments: /


Oil:

Kitty Oil Quantity (# of Quarts in storage): the previous commander told us not to check the oil quantity of Kitty because we have to turn the all generator off!

Honey Oil Quantity (# of Quarts in storage): /

ATV Oil Quantity (# of Quarts in storage): full

Car Oil Quantity (# of Quarts in storage): /


Water Status:

Outside Potable Water Tank Level (inches from bottom): 14,5

Trailer Potable Water Tank Level (inches from bottom): 9,5

Hab Potable Water Tank Level (inches from bottom): 10

Potable Water Meter Reading: 45222,8 (it seems there was a mistake with the value of yesterday)

In to GreenHab Meter Reading: /

Notes/Comments: /


GreenHab:

Condition of plants in Tank 1: /

Condition of plants in Tank 2: /

Condition of plants in Tank 3: /

Greenhab Notes/Comments: One of the two ventilators is on, I don't find the electric source for the other. The thermostat is at 50°F and the temperature inside is 97°F


Observatory:

Telescope Used? (Yes/No): no

Observatory Notes/Comments: there's no telescope


Transportation:

Hab Car (NOT RENTAL)

Used (yes/no): no

Oil Checked (yes/no): no

Tire Status: ok

Odometer at end of day: 124250,5

Notes/Comments on Hab Car: /


ATV 1 (Opportunity):

Used: yes

Oil Checked: no

Fuel consumed : 1/5 tank approximately

Tire Status: ok

Comments: /


ATV 2 (Spirit):

Used: yes

Oil Checked: no

Fuel Consumed: 1/5 tank approximately

Tire Status: ok

Comments: /


ATV 3 (Viking I):

Used: yes

Oil Checked: no

Fuel Consumed: 1/5 tank approximately

Tire Status: ok

Comments: /


ATV 4 (Viking II):

Used: no

Oil Checked: no

Fuel Consumed: 0

Tire Status:

Comments: does not work


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 60

Actual temperature upstairs (Farenheit): 76

Thermostat setting downstairs (Farenheit): /

Actual temperature downstairs (Farenheit): 75


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: internet is still running well


General Engineering Notes/Comments: /


Science Report (Balloon Experiment)

Jonas Offergeld Reporting

Today during the afternoon we started a new experience. We try to build a balloon with a dark plastic sheeting to take aerial pictures.

Crew 94,which was here just a year ago, had started an experience with some balloons filled with helium.

We plan to do the same with a self-made hot-air balloon.

So, we solder a rectangular dark plastic sheeting with a ironer to make cylinder. Then we solder a scare a the top to close the cylinder. finally, we have a cylinder 5 ft diameter and 14 ft high.

We use a plastic tube to make a rigid structure for the base of the balloon.

Today we just constructed the basic structure and tomorrow, when the wind will be weaker, we will try to fill it with hot air outside, thanks to our Bunsen burner.

We hope to fix a camera to the balloon to take some picture of the station.


EVA 8 Report

Thibault Windels Reporting

Time of Departure: 16h10

Time of Arrival: 18h00

Crew Members: S. Viroux, T. Windels

Site Location: 12S 0519861 / 4251222

12S 0520165 / 4248950

12S 0520182 / 4248937

~12S 0515796 / 4251246~

Transit Mode: By ATV

Objectives: The purpose of this EVA was to follow the path of yesterday's crew (without any indications, only with the GPS) in order to take samples for the permeability experiment to the Hab.

Results: We didn't succeed in finding the last location

Lessons Learned: Improve orientation in the field


Chef's Report

Gregory Van Roy Reporting

1a) Today was

A cooking day

1b) Was today a special day celebrated at the main meal with special food

or activity?

Nothing special today

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

Cup a soup Chicken Noodles, Bacon full cooked, Potatoes, Mushrooms and Strawberry drink mix

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

Bacon : 3 packs left

Potatoes : 1/5 of the pack left

Mushroom : 3/4 of the pack left

4) Recipes for anything more complicated than following package

directions exactly, or rehydrating.

None

5) Lessons learned (if any)

None

6) Comments/questions for Kim and Jean, the food study investigators

We don't have dry yeast for bread-machine anymore...